OpenStudy (anonymous):

what happened to the capacitance of a capacitor when we place a conductor between them

OpenStudy (anonymous):

between the plates of the capacitor or what?

OpenStudy (anonymous):

Is the conductor connected to the plates? Or just in the middle?

OpenStudy (anonymous):

it will act as a resistance or any other conductor

OpenStudy (aum):

If the conductor touches both plates of the capacitor, then the capacitor cannot hold any charge and so the effective capacitance will drop to zero. C = Q/V = 0/V = 0.